Fermi National Laboratory
 

Operation Summary for May 26, 2010 - May 28, 2010

 

Wednesday May 26

The day shift began with Operations monitoring the Tevatron (TeV) store 7836 (81.97E30) and the Recycler stash (194.9E10), with the Antiproton Source (Pbar) stacking (20.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, with beam to the Fermilab Test Beam Facility (FTBF), and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, MiniBooNE was off from 9 AM to 2:57 PM due to bad air conditioning units at MI-12. Operations terminated store 7836 at 10:34 AM. Operations prepared CDF and D0 for access and for TeV sector B1 wet engine repairs. Operations began the TeV turn-on and checkout at 2 PM.

During the evening shift, Operations put the TeV into HEP shot setup at 4:37 PM and established store 7839 at 5:56 PM with an average initial luminosity (AIL) of 375.6E30. Poor proton and antiproton beam positions resulted in a quench at D4 at 6:17 PM. Operations recovered the TeV, checked the orbits, and put the TeV into shot setup.

Thursday May 27

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (5.5E10), with Pbar stacking (39.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 7844 at 1:56 AM with an AIL of 193.8E30.

During the day shift, Operations stopped stacking at 10:40 AM for work on the Pbar cryo system. Operations terminated store 7844 at 3 PM. TeV studies followed.

During the evening shift, Operations put the TeV into shot setup at 8:44 PM. Unfortunately, the shot setup ended in a quench at D0 due to instability of the colliding beams.

Friday May 28

The midnight shift began with Operations monitoring the stash (0.9E10), with the Pbar cryo system in cooldown mode, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Pbar resumed stacking at 4:28 AM as their cryo system cooldown would allow. Operations halted the Booster beam at 7 AM in preparation for an access. Operators continued the TeV quench investigation and recovery, and conducted proton studies throughout the shift.

The Plans for Friday and the Weekend

The plans for today are to let Booster personnel make RF repairs in the tunnel, and for TeV personnel to continue their studies using wet squeezes. The Run Coordinator would like to resume stacking, stashing, and storing later this afternoon and continue throughout the weekend.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/28/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory